# Forecasting and Manipulating the Forecasts of Others

### 18 Mar 2026 v2

@@ B.6 Risk-sensitive extension: exponential utility FOC
Appendix C Well-posedness of the deterministic impulse-response fixed point
@@
C.1 Standing assumptions
Assumption C.1.
@@ C.1 Standing assumptions
C.2 The kernel state space and its dynamics
@@
$\mathcal{Z}_{t}:=\bigl(\bar{X}_{t},\;X_{t}(\cdot),\;\{F_{t}^{j}\}_{j=1}^{n}\bigr)\;\in\;\mathbb{R}^{d}\times L^{2}([0,t];\mathbb{R}^{d\times(n+1)d})\times\prod_{j=1}^{n}L^{2}_{\mathrm{sym}}([0,t]^{2};\mathbb{R}^{(n+1)d\times(n+1)d}).$ (C.1)
@@
$\dot{\bar{X}}(t)$ $=A(t)\bar{X}_{t}+\textstyle\sum_{k}\bar{D}^{k}_{t},$ (C.3)
$\dot{X}_{t}(s)$ $=A(t)X_{t}(s)+\textstyle\sum_{k}\mathcal{D}^{k}_{t}(s),\qquad X(s,s)=\Sigma(s),$ (C.4)
$\partial_{t}F_{t}^{j}(u,s)$ $=\widetilde{X}^{j}_{t}(u)^{\top}P^{j}(t)\,\widetilde{X}^{j}_{t}(s),\qquad F_{0}^{j}=0,$ (C.5)
with boundary $F_{t}^{j}(u,t)=\widetilde{X}^{j}_{t}(u)^{\top}\sqrt{P^{j}(t)}\,E^{j}$, and primitive-noise control $\mathcal{D}^{k}_{t}(s):=D^{k}_{t}(s)\Pi^{k}+\int_{0}^{t}D^{k}_{t}(u)\,F_{t}^{k}(u,s)\,du$.
The policy $\mathbf{D}$ enters (C.3)–(C.4) but does not enter (C.5) directly: $F^{j}$ depends on $\mathbf{D}$ only through $X$ via (C.2).
C.3 The Hamilton–Jacobi–Bellman equation
Let $V^{i}(t,\mathcal{Z})$ denote the equilibrium continuation cost for player $i$ from kernel state $\mathcal{Z}$ at time $t$. Since the kernel-state dynamics are deterministic, the HJB is a first-order terminal-value problem:
$\partial_{t}V^{i}+D_{\bar{X}}V^{i}\cdot\dot{\bar{X}}+\langle D_{X}V^{i},\,\dot{X}\rangle+\sum_{j=1}^{n}\langle D_{F^{j}}V^{i},\,\dot{F}^{j}\rangle+\ell^{i}(t,\mathcal{Z},\mathbf{D})=0,\qquad V^{i}(T,\cdot)=g^{i}(\cdot),$ (C.6)
where $\langle\cdot,\cdot\rangle$ denotes the $L^{2}$ pairing, $\ell^{i}$ is the running cost, and $g^{i}$ is the terminal cost, both evaluated at the kernel state.
Player $i$’s control $(\bar{D}^{i}_{t},D^{i}_{t}(\cdot))$ enters $\dot{\bar{X}}$ and $\dot{X}$ linearly but does not enter $\dot{F}^{j}$; the Hamiltonian minimization over $D^{i}$ therefore involves only $D_{\bar{X}}V^{i}$, $D_{X}V^{i}$, and the quadratic effort cost, giving the first-order condition
$\bar{D}^{i}_{t}=-(G^{DD,i}(t))^{-1}D_{\bar{X}}V^{i},\qquad D^{i}_{t}(u)=-(G^{DD,i}(t))^{-1}D_{X}V^{i}(u).$ (C.7)
Along the equilibrium path, $D_{\bar{X}}V^{i}=\bar{H}^{X,i}_{t}$ and $D_{X}V^{i}(\cdot)=H^{X,i}_{t}(\cdot)$, recovering Corollary 3.5.
The information wedge as a shadow price.
The $\sum_{j}\langle D_{F^{j}}V^{i},\dot{F}^{j}\rangle$ term in (C.6) captures the indirect effect of the kernel state’s evolution on player $i$’s continuation cost through opponents’ filters. Along the equilibrium path, differentiating through the chain $D^{i}\to X\to\widetilde{X}^{k}\to F^{k}$ identifies
$\bar{\mathcal{V}}^{i}_{t}=\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz=\sum_{k\neq i}\langle D_{F^{k}}V^{i},\,\partial_{X}\dot{F}^{k}\rangle,$ (C.8)
i.e. the mean information wedge of Theorem 3.2 is the shadow price of the opponent’s filter evolution, contracted against the sensitivity of that evolution to the state kernel. The kernel wedge $\mathcal{V}^{i}_{t}(r)$ admits the same interpretation with $\bar{H}^{k,i}$ replaced by $H^{k,i}(\cdot,\cdot,r)$.
Endogenous precision.
If $P^{i}(\cdot)$ is player $i$’s control subject to attention cost $\int_{0}^{T}c^{i}(t,P^{i}(t))\,dt$, then $P^{i}$ enters (C.6) only through $\dot{F}^{i}$ via (C.5), and the FOC
$\bigl\langle D_{F^{i}}V^{i},\;\widetilde{X}^{i}_{t}(\cdot)^{\top}(\cdot)\,\widetilde{X}^{i}_{t}(\cdot)\bigr\rangle=\frac{\partial c^{i}}{\partial P^{i}}(t,P^{i}(t))$ (C.9)
decouples from the FOC for $D^{i}$: the left side is the marginal value of accelerating the filter, which vanishes as $\mathcal{P}_{t}^{i}\to I$ (Lemma C.3), guaranteeing an interior solution when $\partial c^{i}/\partial P^{i}>0$.
C.4 A priori bounds from projection structure
@@
Lemma C.2 (Projection bounds).
@@
Equation (C.5) adds a positive semidefinite increment to $F_{t}^{j}$, so $\mathcal{P}_{t}^{j}$ is non-decreasing. Since $\widetilde{X}^{j}(t,\cdot)=(I-\mathcal{P}_{t}^{j})X(t,\cdot)$ by (C.2), Lemma C.2(ii) gives the bound. ∎
These bounds hold uniformly in the precision paths, the policy profile, and the horizon $T$. The quadratic right-hand side of (C.5) is self-limiting: as $\mathcal{P}_{t}^{j}$ grows toward $I$, the complementary projection contracts, driving $\widetilde{X}^{j}\to 0$ and decelerating $F^{j}$. This prevents finite-time blowup despite the quadratic nonlinearity and is the primitive noise-linear manifestation of Kalman variance collapse.
C.5 Well-posedness
Definition C.1.
@@ C.5 Well-posedness
Definition C.2 (Best-response operator).
@@
The following estimates are proved by Grönwall arguments on the forward system (C.3)–(C.5) (using the a priori bounds of Lemmas C.2C.3 to control the quadratic nonlinearity in (C.5)) and by backward Grönwall on the linear adjoint system (B.11)–(B.14) (using Cauchy–Schwarz on the belief-adjoint coupling integrals $\int_{0}^{t}P^{k}\widetilde{X}^{k}\,\bar{H}^{k,i}\,dz$).
Proposition C.4 (Bounds and Lipschitz continuity).
Under Assumption C.1, for $\mathbf{D}\in\mathbb{B}_{T}^{M}$:
@@
For the forward system: $(\bar{X},X)$ solve linear equations with bounded coefficients once $\mathcal{D}^{k}$ is controlled. Since $\mathcal{D}^{k}$ involves $F^{k}$ through a linear integral, and $\|\mathcal{F}_{t}^{k}\|_{\mathrm{op}}\leq 2$ (Lemma C.2), the coupled system for $\phi(t):=\|X\|_{L^{\infty}(\Delta_{t})}$ closes as $\phi(t)\leq C(M)+C(M)\int_{0}^{t}\phi(r)\,dr$.
For the Lipschitz bound: the variation $\delta\mathcal{D}^{k}$ decomposes into a direct term bounded by $C(M)\|\delta\mathbf{D}\|$ and an indirect term through $\delta F^{k}$, which couples to $\delta\widetilde{X}^{k}$ and back to $\delta X$ via (C.2). A Grönwall argument on the aggregate $\Psi(t):=\|\delta X\|_{L^{\infty}(\Delta_{t})}+\max_{j}\sup_{r\leq t}\|\delta\widetilde{X}^{j}(r,\cdot)\|_{L^{2}}$ gives $\Psi(T)\leq C(M)Te^{C(M)T}\|\delta\mathbf{D}\|$; the factor of $T$ arises because the policy enters through integrals of length $\leq T$.
@@
Theorem C.5 (Short-horizon equilibrium).
Under Assumption C.1, there exists $T^{*}>0$ depending only on $(n,d,L,\lambda)$ such that for $\in(0,T^{*}], 𝒯\mathcal{T}$ is a contraction on $\mathbb{B}_{T}^{M_{0}}$ for a suitable $M_{0}$. In particular, there exists a unique Nash equilibrium in noise-state linear strategies with bounded kernels, Picard iterates converge geometrically, and the equilibrium is Nash over the full admissible $L^{2}$ class.
@@
By Proposition C.4(iii), $\mathrm{Lip}(\mathcal{T})\leq\lambda^{-1}C(M_{0})Te^{C(M_{0})T}$. Self-mapping ($\mathcal{T}:\mathbb{B}_{T}^{M_{0}}\to\mathbb{B}_{T}^{M_{0}}$) holds because at $M=0$ the belief-adjoint decouples ($D^{k}\equiv 0$ kills the forcing in (B.12)), giving finite $K_{\mathrm{bwd}}(0,T)$; the intermediate value theorem yields $M_{0}$. Since $M_{0}$ is bounded as $T\to 0$, choosing $T^{*}$ small enough makes $\mathrm{Lip}(\mathcal{T})\leq\tfrac{1}{2}$. Banach’s theorem gives existence, uniqueness, and geometric convergence. Corollary 3.5 extends to the full $L^{2}$ class. ∎
Theorem C.6 (Global existence).
Under Assumption C.1, for every $T>0$ there exists a Nash equilibrium over the full admissible $L^{2}$ class.
@@
Apply the Schauder fixed-point theorem to $\mathcal{T}$ on $\mathbb{B}_{T}^{M_{0}}$, working in the topology $\tau:=(C([0,T])\times L^{2}(\Delta_{T}))^{n}$. Self-mapping holds by the same argument as above. $\tau$-closedness of $\mathbb{B}_{T}^{M_{0}}$: weak-* lower semicontinuity of the $L^{\infty}$ norm preserves the bound under $L^{2}$ limits. $\tau$-continuity of $\mathcal{T}$: on the bounded set $\mathbb{B}_{T}^{M_{0}}$, $L^{2}$-convergence of policy kernels propagates through the forward linear equations (via Cauchy–Schwarz) and the linear backward system. $\tau$-precompactness of the image: the uniform time-derivative bound on $\bar{H}^{X,i}$ (Proposition C.4(ii)) gives equicontinuity of $\bar{D}^{i,\mathrm{new}}$ in $C([0,T])$; the same bound gives pointwise-in-$u$ equicontinuity of $D^{i,\mathrm{new}}(t,u)$ in $t$, which with the uniform $L^{\infty}$ bound yields $L^{2}$-equicontinuity of $t\mapsto D^{i,\mathrm{new}}(t,\cdot)$ and hence sequential compactness in $L^{2}(\Delta_{T})$ by the vector-valued Arzelà–Ascoli theorem ([1, Theorem 1.3.2]). ∎
Remark C.1 (Uniqueness beyond the contraction regime).
@@
Theorem C.6 gives existence but not uniqueness for $T>T^{*}$. The dissipation of Lemma C.3 suggests enough structure to preclude multiplicity, but a proof requires monotonicity estimates beyond the Grönwall bounds used here. In Section 4, Picard iteration converges to the same limit from multiple initial conditions for all parameter values tested.
Remark C.2 (Stochastic coefficients).
@@
If $A(t,\omega)$, $P^{i}(t,\omega)$, or the cost matrices depend on a public factor process $\xi_{t}$ (e.g. a finite-state Markov chain representing regime switches or stochastic automation events), the kernel state augments to $(\mathcal{Z}_{t},\xi_{t})$. The projection bounds of Lemmas C.2C.3 are algebraic properties of orthogonal projections and hold pathwise, so the forward well-posedness extends to each sample path between jumps of $\xi$. Each jump triggers a regime transition in which filtering gains, strategies, and the information wedge shift discontinuously, while $\mathcal{Z}$ is continuous across jumps. The backward ODE is replaced by a backward stochastic differential equation in the adjoint kernels, driven by jumps of $\xi$; the deterministic Grönwall arguments extend to pathwise bounds conditional on the regime path.
Appendix D Kyle–Back derivations
This appendix derives the deviation propagation system, the first-order condition (5.3), and the closed backward ODE (5.4) stated in Section 5.
D.1 Drift kernels
The market maker’s drift kernel is $C^{0}(t,s):=\sigma_{0}^{-1}\mathcal{D}_{\Sigma,t}(s)$; for trader $k$:
$C^{k}(t,s):=\begin{pmatrix}\sigma_{0}^{-1}\mathcal{D}_{\Sigma,t}(s)\\[2.0pt] \sqrt{p_{k}}\,\sigma_{V}\,E^{V}\end{pmatrix}\!,\qquad\Pi_{\mathcal{E}}^{k}:=\Pi^{0}+\Pi^{k}.$ (D.1)
The unresolved kernels $\widetilde{C}^{0}$ and $\widetilde{C}^{k}$ are defined by the projection decomposition (A.12) with $C^{j}$ replacing $\sqrt{P^{j}}X$ and $\Pi_{\mathcal{E}}^{j}$ replacing $\Pi^{j}$.
D.2 Deviation propagation

### 23 Mar 2026 v3

@@ C.2 Two channels from a precision perturbation
Appendix D Well-posedness of the deterministic impulse-response fixed point
@@
D.1 Standing assumptions
Assumption D.1.
@@ D.1 Standing assumptions
D.2 The kernel state space and its dynamics
@@
$\mathcal{Z}_{t}:=\bigl(\bar{X}_{t},\;X_{t}(\cdot),\;\{F_{t}^{j}\}_{j=1}^{n}\bigr)\;\in\;\mathbb{R}^{d}\times L^{2}([0,t];\mathbb{R}^{d\times(n+1)d})\times\prod_{j=1}^{n}L^{2}_{\mathrm{sym}}([0,t]^{2};\mathbb{R}^{(n+1)d\times(n+1)d}).$ (D.1)
@@
$\dot{\bar{X}}(t)$ $=A(t)\bar{X}_{t}+\textstyle\sum_{k}\bar{D}^{k}_{t},$ (D.3)
$\dot{X}_{t}(s)$ $=A(t)X_{t}(s)+\textstyle\sum_{k}\mathcal{D}^{k}_{t}(s),\qquad X(s,s)=\Sigma(s),$ (D.4)
$\partial_{t}F_{t}^{j}(u,s)$ $=\widetilde{X}^{j}_{t}(u)^{\top}P^{j}(t)\,\widetilde{X}^{j}_{t}(s),\qquad F_{0}^{j}=0,$ (D.5)
with boundary $F_{t}^{j}(u,t)=\widetilde{X}^{j}_{t}(u)^{\top}\Gamma^{j}(t)\,E^{j}$, and primitive-noise control $\mathcal{D}^{k}_{t}(s):=D^{k}_{t}(s)\Pi^{k}+\int_{0}^{t}D^{k}_{t}(u)\,F_{t}^{k}(u,s)\,du$.
The policy $\mathbf{D}$ enters (D.3)–(D.4) but does not enter (D.5) directly: $F^{j}$ depends on $\mathbf{D}$ only through $X$ via (D.2).
D.3 The Hamilton–Jacobi–Bellman equation
Since the kernel-state dynamics (D.3)–(D.5) are deterministic, the equilibrium continuation cost $V^{i}(t,\mathcal{Z})$ satisfies a first-order terminal-value PDE (no diffusion term):
$\partial_{t}V^{i}+D_{\bar{X}}V^{i}\cdot\dot{\bar{X}}+\langle D_{X}V^{i},\,\dot{X}\rangle+\sum_{j=1}^{n}\langle D_{F^{j}}V^{i},\,\dot{F}^{j}\rangle+\ell^{i}(t,\mathcal{Z},\mathbf{D})=0,\qquad V^{i}(T,\cdot)=g^{i}(\cdot),$ (D.6)
where $\langle\cdot,\cdot\rangle$ denotes the $L^{2}$ pairing. Player $i$’s control enters $\dot{\bar{X}}$ and $\dot{X}$ but not $\dot{F}^{j}$, so the Hamiltonian minimization yields
$\bar{D}^{i}_{t}=-(G^{DD,i}(t))^{-1}D_{\bar{X}}V^{i},\qquad D^{i}_{t}(u)=-(G^{DD,i}(t))^{-1}D_{X}V^{i}(u),$ (D.7)
recovering Corollary 4.5 along the equilibrium path. The remaining $\sum_{j}\langle D_{F^{j}}V^{i},\dot{F}^{j}\rangle$ term captures the indirect cost of opponents’ filter evolution; differentiating through $D^{i}\to X\to\widetilde{X}^{k}\to F^{k}$ identifies the mean information wedge as a shadow price:
$\bar{\mathcal{V}}^{i}_{t}=\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz=\sum_{k\neq i}\langle D_{F^{k}}V^{i},\,\partial_{X}\dot{F}^{k}\rangle.$ (D.8)
The kernel wedge $\mathcal{V}^{i}_{t}(r)$ admits the same interpretation with $\bar{H}^{k,i}$ replaced by $H^{k,i}(\cdot,\cdot,r)$.
D.4 A priori bounds from projection structure
@@
Lemma D.2 (Projection bounds).
@@
Equation (D.5) adds a positive semidefinite increment to $F_{t}^{j}$, so $\mathcal{P}_{t}^{j}$ is non-decreasing. Since $\widetilde{X}^{j}(t,\cdot)=(I-\mathcal{P}_{t}^{j})X(t,\cdot)$ by (D.2), Lemma D.2(ii) gives the bound. ∎
These bounds hold uniformly in the precision paths, the policy profile, and the horizon $T$. The quadratic right-hand side of (D.5) is self-limiting: as $\mathcal{P}_{t}^{j}$ grows toward $I$, the complementary projection contracts, driving $\widetilde{X}^{j}\to 0$ and decelerating $F^{j}$. This prevents finite-time blowup despite the quadratic nonlinearity—the noise-state analogue of Kalman variance collapse.
D.5 Well-posedness
Definition D.1.
@@ D.5 Well-posedness
Definition D.2 (Best-response operator).
@@
The following estimates are proved by Grönwall arguments on the forward system (D.3)–(D.5) (using the a priori bounds of Lemmas D.2D.3 to control the quadratic nonlinearity in (D.5)) and by backward Grönwall on the linear adjoint system (B.11)–(B.14) (using Cauchy–Schwarz on the belief-adjoint coupling integrals $\int_{0}^{t}P^{k}\widetilde{X}^{k}\,\bar{H}^{k,i}\,dz$).
Proposition D.4 (Bounds and Lipschitz continuity).
Under Assumption D.1, for $\mathbf{D}\in\mathbb{B}_{T}^{M}$:
@@
For the forward system: $(\bar{X},X)$ solve linear equations with bounded coefficients once $\mathcal{D}^{k}$ is controlled. Since $\mathcal{D}^{k}$ involves $F^{k}$ through a linear integral, and $\|\mathcal{F}_{t}^{k}\|_{\mathrm{op}}\leq 2$ (Lemma D.2), the coupled system for $\phi(t):=\|X\|_{L^{\infty}(\Delta_{t})}$ closes as $\phi(t)\leq C(M)+C(M)\int_{0}^{t}\phi(r)\,dr$.
For the Lipschitz bound: the variation $\delta\mathcal{D}^{k}$ decomposes into a direct term bounded by $C(M)\|\delta\mathbf{D}\|$ and an indirect term through $\delta F^{k}$, which couples to $\delta\widetilde{X}^{k}$ and back to $\delta X$ via (D.2). A Grönwall argument on the aggregate $\Psi(t):=\|\delta X\|_{L^{\infty}(\Delta_{t})}+\max_{j}\sup_{r\leq t}\|\delta\widetilde{X}^{j}(r,\cdot)\|_{L^{2}}$ gives $\Psi(T)\leq C(M)Te^{C(M)T}\|\delta\mathbf{D}\|$; the factor of $T$ arises because the policy enters through integrals of length $\leq T$.
@@
Theorem D.5 (Short-horizon equilibrium).
Under Assumption D.1, there exists $T^{*}>0$ depending only on $(n,d,L,\lambda)$ such that for $\in(0,T^{*}], 𝒯\mathcal{T}$ is a contraction on $\mathbb{B}_{T}^{M_{0}}$ for a suitable $M_{0}$. In particular, there exists a unique Nash equilibrium in noise-state linear strategies with bounded kernels, Picard iterates converge geometrically, and the equilibrium is Nash over the full admissible $L^{2}$ class.
@@
By Proposition D.4(iii), $\mathrm{Lip}(\mathcal{T})\leq\lambda^{-1}C(M_{0})Te^{C(M_{0})T}$. Self-mapping ($\mathcal{T}:\mathbb{B}_{T}^{M_{0}}\to\mathbb{B}_{T}^{M_{0}}$) holds because at $M=0$ the belief-adjoint decouples ($D^{k}\equiv 0$ kills the forcing in (B.12)), giving finite $K_{\mathrm{bwd}}(0,T)$; the intermediate value theorem yields $M_{0}$. Since $M_{0}$ is bounded as $T\to 0$, choosing $T^{*}$ small enough makes $\mathrm{Lip}(\mathcal{T})\leq\tfrac{1}{2}$. Banach’s theorem gives existence, uniqueness, and geometric convergence. Corollary 4.5 extends to the full $L^{2}$ class. ∎
Theorem D.6 (Global existence).
Under Assumption D.1, for every $T>0$ there exists a Nash equilibrium over the full admissible $L^{2}$ class.
@@
Apply the Schauder fixed-point theorem to $\mathcal{T}$ on $\mathbb{B}_{T}^{M_{0}}$, working in the topology $\tau:=(C([0,T])\times L^{2}(\Delta_{T}))^{n}$. Self-mapping holds by the same argument as above. $\tau$-closedness of $\mathbb{B}_{T}^{M_{0}}$: weak-* lower semicontinuity of the $L^{\infty}$ norm preserves the bound under $L^{2}$ limits. $\tau$-continuity of $\mathcal{T}$: on the bounded set $\mathbb{B}_{T}^{M_{0}}$, $L^{2}$-convergence of policy kernels propagates through the forward linear equations (via Cauchy–Schwarz) and the linear backward system. $\tau$-precompactness of the image: the uniform time-derivative bound on $\bar{H}^{X,i}$ (Proposition D.4(ii)) gives equicontinuity of $\bar{D}^{i,\mathrm{new}}$ in $C([0,T])$; the same bound gives pointwise-in-$u$ equicontinuity of $D^{i,\mathrm{new}}(t,u)$ in $t$, which with the uniform $L^{\infty}$ bound yields $L^{2}$-equicontinuity of $t\mapsto D^{i,\mathrm{new}}(t,\cdot)$ and hence sequential compactness in $L^{2}(\Delta_{T})$ by the vector-valued Arzelà–Ascoli theorem ([2, Theorem 1.3.2]). ∎
Remark D.1 (Uniqueness beyond the contraction regime).
@@
Theorem D.6 gives existence but not uniqueness for $T>T^{*}$. The dissipation of Lemma D.3 suggests enough structure to preclude multiplicity, but a proof requires monotonicity estimates beyond the Grönwall bounds used here. In Section 2, Picard iteration converges to the same limit from multiple initial conditions for all parameter values tested.
Remark D.2 (Stochastic coefficients).
@@
If primitives depend on a public factor process $\xi_{t}$ (e.g. a finite-state Markov chain representing regime switches), the kernel state augments to $(\mathcal{Z}_{t},\xi_{t})$. The projection bounds of Lemmas D.2D.3 are algebraic and hold pathwise, so forward well-posedness extends between jumps of $\xi$. Each jump triggers a regime transition in which filtering gains, strategies, and the information wedge shift discontinuously, while $\mathcal{Z}$ is continuous across jumps. The backward ODE becomes a BSDE driven by jumps of $\xi$; the Grönwall arguments extend to pathwise bounds conditional on the regime path.
Appendix E Kyle–Back derivations
This appendix derives the price-impact kernel $l(\tau,t)$, the first-order condition (5.3), and the backward ODE (5.4), with all opponents’ noise-state linear strategy maps held fixed.
E.1 The price-impact kernel and Kyle’s lambda